### v2

# DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

Abstract
Native multimodal large language models (MLLMs) restructure a single large language model (LLM) into a spoken language model (SLM) capable of both speech and text generation. Compared to modular and aligned MLLMs, native MLLMs preserve richer paralinguistic features such as emotion and prosody, and generate speech responses directly within the backbone LLM rather than using a separate speech decoder. This integration also results in lower response latency and smoother interaction. However, native MLLMs suffer from catastrophic forgetting and performance degradation because the available paired speech-text data is insufficient to support the pretraining of MLLMs compared to the vast amount of text data required to pretrain text LLMs. To address this issue, we propose DeepTalk, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs. Meanwhile, the end-to-end dialogue latency remains within $0.5$ seconds, ensuring a seamless and intelligent speech interaction experience. Code and models are released at https://github.com/talkking/DeepTalk.
1 Introduction
Since GPT-4o GPT4o has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models. As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) radford2023robust ; bai2024seed ; povey2011kaldi , the resulting text is then comprehended by a large language model (LLM) liu2024deepseek ; yang2024qwen2 ; glm2024chatglm to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) du2024cosyvoice ; ren2019fastspeech ; wang2025spark model. However, this cascaded structure results in high latency, and errors from each module can accumulate and propagate, leading to degraded performance. Therefore, recent speech LLMs adopt an end-to-end approach, using a unified model to process both speech and text simultaneously, which enhances the capabilities of speech understanding and generation. At the same time, this approach significantly reduces interaction latency and provides a more natural and smooth voice interaction experience.
Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal chen2025minmo . Representatives of the former include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , Freeze-Omni wang2024freeze , and VITA-1.5 fu2025vita . These models connect a speech encoder and a speech decoder to the LLM to handle audio understanding and audio generation tasks separately, placing semantic understanding within the LLM module. This maximally the retains LLM’s general capabilities. Furthermore, since the LLM itself does not require retraining, much less speech-text paired pretraining data is needed.
Representative of the latter including Mini-Omni xie2024mini , GLM-4-Voice zeng2024glm , Moshi defossez2024moshi , and LUCY gao2025lucy use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens. Compared to modular aligned multimodal approaches, native ones greatly reduces the latency of speech token output. More paralinguistic information in speech, such as emotion and prosody, is retained within the SLM rather than in the speech decoder, benefiting the full expression of paralinguistic information. However, since native multimodal models expand the original LLM’s vocabulary and require retraining, they need a large amount of audio-text paired pretraining data. Therefore, in real-world scenarios where audio data is much less abundant than text data, native multimodal models are prone to catastrophic forgetting, severely impairing the original LLM’s language capabilities.
In this paper, to alleviate the catastrophic forgetting faced by native multimodal models zhai2023investigating , we propose DeepTalk, a model that applies the mixture of experts (MoE) for multimodal learning li2025uni ; li2024cumo ; zhong2024moextend ; shen2024mome to effectively isolate modality-specific knowledge. Since most speech dialogue data are conversational and colloquial, while text LLMs are primarily trained on formal written text, training LLMs with colloquial speech dialogue data can lead to less standardized written outputs, resulting in decreased language performance on certain LLM benchmarks. Therefore, we assign different parameters within the LLM to specialize in instruction data from different modalities, allowing modality experts to focus on single-modality training and avoid interference between modalities. Finally, we use cross-modal instruction data to jointly train the modality experts, enabling the model to output both speech and text simultaneously.
To minimize the damage to the original text LLM, we propose an adaptive modality expert selection strategy, which dynamically selects modality experts based on the MoE model’s modality load on different data. Experts with a high speech token load but a low text token load are selected as audio experts. After several iterations of modality expert selection and training, DeepTalk achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
In summary, the contributions of this paper are as follows:
  • We propose DeepTalk, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge. At the same time, modalities can collaborate to jointly output multimodal results, enabling both colloquial and formal language to coexist within a single model.
  • We introduce an adaptive modality expert selection method based on modality load, which further reduces the damage to the original text LLM and alleviates the problem of catastrophic forgetting.
  • To the best of our knowledge, this is the first native multimodal large speech interaction model based on the MoE architecture. It opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
Figure 1: Comparison of DeepTalk and existing voice interaction systems. Compared to modular MLLMs, DeepTalk embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM. Through adaptive modality expert selection and a three-stage training process for modality experts, DeepTalk effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
2 Related Work
2.1 End-to-End Speech Interaction System
Speech is among the most important signal in human communication, as it not only conveys semantic content but also carries various paralinguistic information such as emotion, speech rate, intonation, and timbre. As a result, speech interaction has become an increasingly important mode of human-computer interaction. Traditional speech interaction systems are typically composed of a pipeline of separately optimized ASR, LLM, and TTS modules. Such systems not only suffer from high interaction latency, but also accumulate errors from each module, leading to a less satisfactory overall user experience. Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach ji2024wavchat .
Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM kong2020panns ; chu2024qwen2 ; chu2023qwen ; das2024speechverse . This approach not only provides the LLM with richer speech information than using pure text input, but also reduces the latency caused by ASR recognition. However, these models still require an additional TTS module to achieve end-to-end speech interaction. To provide rich speech information, LLaMA-Omni fang2024llama feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens. Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss graves2006connectionist is used to align the upsampled hidden states with the audio token outputs. Meanwhile, to reduce latency for real-time speech interaction, a predefined chunk size is set to enable the vocoder to synthesize speech in a streaming manner. Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of $82$Hz and adopts a delayed decoding strategy copet2023simple ; peng2024voicecraft ; lyth2024natural to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference. Minmo chen2025minmo integrates a language model with an autoregressive CosyVoice2 decoder du2024cosyvoice . The hidden states of the language model are fed into the CosyVoice2 decoder to autoregressively predict speech tokens, which are then passed to a vocoder for waveform synthesis.
Fully end-to-end approaches integrate the TTS module into the LLM, allowing the LLM itself to generate both text and speech tokens, rather than relying on a separate speech decoder or TTS module for speech synthesis. Discrete speech tokens are typically generated by neural audio codecs defossez2022high ; ji2024wavtokenizer ; zhang2023speechtokenizer or self-supervised models lakhotia2021generative ; hsu2021hubert ; chen2022wavlm . Currently, there are two mainstream modeling paradigm for end-to-end speech systems. The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice zeng2024glm and Spirit-LM nguyen2025spirit , where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens. The second is parallel audio-text modeling, represented by models such as Mini-Omni xie2024mini , Moshi defossez2024moshi , Slam-Omni chen2024slam and LUCY gao2025lucy . In this paradigm, text and audio tokens can be generated in parallel: while generating text tokens, multiple LM heads are used to simultaneously predict multiple audio tokens. Compared to the interleaved approach, parallel modeling compresses speech tokens into shorter sequences, enabling the modeling of higher-bitrate speech tokens.
Modular Speech Language Models  Modular large speech models are primarily built on an architecture where a LLM is directly connected to an audio encoder and decoder through adapters. Examples include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , and Freeze-Omni wang2024freeze , as shown in Fig. 1 b). In this structure, the LLM is only responsible for textual decoding, while speech decoding is handled by the audio decoder. As a result, the language capabilities of the LLM are largely preserved. Additionally, Minmo employs Low-Rank Adaptation (LoRA) hu2022lora for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities. However, such modular speech-language models face several challenges. For instance, the transmission of speech information relies entirely on the audio decoder, with the LLM only responsible for conveying textual content. This means that the expression of paralinguistic information in speech is completely dependent on the speech decoder, resulting in underutilization of the LLM. Furthermore, there are additional issues such as low deployment efficiency and the complexity of balancing the capacities of the LLM and the SLM.
Native Speech Language Models Native speech language models reconstruct the LLM into an SLM, enabling the LLM to output both text and speech tokens simultaneously, as shown in Fig. 1 c). Currently, most mainstream native multimodal approaches such as Mini-Omni xie2024mini , Moshi defossez2024moshi , GLM-4-Voice zeng2024glm , and LUCY gao2025lucy have the LLM handle both text and speech decoding. This approach offers low latency, easy deployment, and allows a unified model to process both audio and text decoding, which better aligns with end-to-end requirements. However, without large-scale speech-text paired data, retraining the LLM in this way can lead to catastrophic forgetting and performance degradation.
2.2 Multimodal Mixture-of-Experts
BEiT-3 wang2023image , VLMo bao2022vlmo , and Uni-MoE li2025uni employ specific modality expert groups to capture modality-specific information. MoExtend zhong2024moextend further expands modality experts horizontally on top of existing LLMs to capture additional modality information. CuMo li2024cumo extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME shen2024mome uses multiple feature encoders as modality expert groups to encode multi-dimensional features. VL-MoE shen2023scaling and MoE-LLaVa lin2024moe introduce mixture-of-experts (MoE) to improve training and deployment. MoMa lin2024moma pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens. Inspired by these works, we introduce multimodal mixture-of-experts (speech experts and text experts) into end-to-end large speech interaction models for native MLLM pretraining. At the same time, we use the modality load of each expert to adaptively select modality experts, a method called Adaptive Modality-Specific MoE, to address the catastrophic forgetting problem aforementioned.
3 Methods
3.1 Architecture
The overall architecture of the model is shown in Fig. 2, which mainly consists of an Audio Encoder, a connector, an MoE backbone, and a streaming codec. Let $X_{i}^{A}$ and $X_{i}^{T}$denote the audio input and text input of the $i$-th sample, respectively, and $Y_{i}^{A}$ and $Y_{i}^{T}$ denote the audio output and text output of the $i$-th sample, respectively. The model’s text and speech responses are represented as $Y^{T}\in V$ and $Y^{A}\in U$, where $V$ and $U$ denote the text vocabulary and the speech codec vocabulary, respectively. The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model, as shown below:
$H_{i}^{A}=Adapter(Encoder(X_{i}^{A})),$ (1)
$H_{i}^{T}=Embedding(X_{i}^{T}),$ (2)
$H_{i}=mean(H_{i}^{A},H_{i}^{T}),$ (3)
where $mean$ here refers to taking the average of the two embeddings. The loss L of the model over N samples can be defined as:
$\textit{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}logP(Y_{i,t}^{T},Y_{i,t}^{A}|Y_{i,<t}^{T},Y_{i,<t}^{A},H_{i}),$ (4)
where $T_{i}$ donates the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
Audio Encoder and Adapter We use Whisper-medium radford2023robust as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads. It includes multiple convolutional downsampling layers that downsample the speech features by a factor of 4. We use an 80-dimensional log Mel-filterbank with a 25ms window length computed every 10ms as the input to the audio encoder. A convolution-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features. Finally, for the input $X_{i}^{A}$, as shown in Equation 1, we obtain the hidden state $H_{i}^{A}$.
Figure 2: The DeepTalk model architecture.
Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better.
Model Seed-TTS $\downarrow$ $\downarrow$ $\downarrow$ Seed-TTS 1.12 2.25 7.59 CosyVoice 3.63 4.29 11.75 CosyVoice2 1.45 2.57 6.83 GLM-4-Voice 2.91 2.10 - VITA-1.5 8.44 2.63 - DeepTalk(topk=10) 1.87 7.34 7.53 DeepTalk(topk=1) 1.68 6.28 7.34 +DPO 1.41 3.25 7.29
Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite liu2024deepseek , as shown in Fig. 2. Each MoE layer is adaptively divided into modality expert groups for specialized training and learning of the corresponding modality. Specifically, the text expert group is responsible for generating written text such as mathematical code, while the speech expert group is responsible for generating spoken language such as emotional speech response. In addition, there is a shared modality expert group for learning common knowledge across both modalities such as daily conversations. The weights among modality expert groups are allocated through modality routing, as follows:
$P(X_{l}^{AT})_{i}=\frac{e^{f(X_{l}^{AT})}}{\sum_{j}^{M+N}e^{f(X_{l}^{AT})_{j}}},$ (5)
$MoE(X_{l}^{AT})=\sum_{i=1}^{L}P(X_{l}^{AT})_{i}\cdot e(X_{l}^{AT})_{i},$ (6)
where $X_{l}^{AT}$ denotes the averaged feature input of speech and text at the $l$-th layer, $L$ represents the total length of the sequence after concatenating speech and text, and $M$ and $N$ denote the number of speech experts and text experts, respectively. $f(x)=W\cdot x$ is the linear modality router to produce expert assignment probabilities and $W\in R^{h\times d}$. $h$ is the last dimension of hidden states and $d$ is the total number of experts.
Audio-Text Modeling Following parallel modeling paradigm defossez2024moshi ; xie2024mini ; chen2024slam ; gao2025lucy , we use different heads to process hidden states, generating both text and audio tokens chen2024slam ; defossez2024moshi in one decoding step. Since the input to the LLM is the average of the text and audio representations instead of native text representations, maintaining the original capabilities of the LLM can be challenging. To enhance models’ text ability, we apply batch parallel decoding xie2024mini ; gao2025lucy , Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response. The generated text tokens from the first sample are of better quality than those from the second and are used to substitute the latter in the second sample.
Audio Codec Decoder We use SNAC codec siuzdak2024snac to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz. Combining speech and text tokens results in eight layers of labels, so we use eight LM heads to simultaneously predict one text token and seven audio token at each decoding step. We follow the delay pattern of MusicGen (copet2024simple, ) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.
3.2 Adaptive Modality-Specific Mixture-of-Experts
The adaptive modality expert selection strategy is primarily designed for an original textual LLM, dynamically distinguishing between audio experts and text experts based on modality load. The goal is to ensure minimal impairment to the language capabilities of the original LLM. Specifically, as shown in Algorithm 1, an initial MoE model that has undergone stage 1 modality alignment training is used. For unimodal input data, the token load rates of all experts for both modalities are calculated, as illustrated in Figure 4. According to step 1 in Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts. As shown in Figure 4, for example, in the first layer, the expert with index 5 (starting from 0) is selected as an audio expert because its load rate for the audio modality is significantly higher than that for the text modality. This adaptive selection approach fully utilizes the text experts with low load in the original MoE LLM, while also ensuring that the selected experts are suitable for processing the audio modality. In this way, the strategy minimizes the impact on the original LLM while fitting the audio modality as much as possible.
Algorithm 1 Adaptive Modality Expert Partitioning and Training
0:  Data with audio input and audio-text output $D^{A}$, Data with text input and audio-text output $D^{T}$, The MoE LLM model $M$ after the first stage of modality alignment.
1:  Initialize $k$ = Number of audio experts, $N$ = Total number of experts.
2:  Step 1. Adaptive partitioning of modality experts
3:   Audio modality token load rate $E^{A}$ = $M(D^{A})$   # Inference with model $M$ on dataset $D^{A}$.
4:   Text modality token load rate $E^{T}$ = $M(D^{T})$   # Inference with model $M$ on dataset $D^{T}$.
5:   _$\_,AE$ = $topk(E^{A}*(1-E^{T}),k)$   # Compute the index $AE$ of the selected audio expert.
6:   _$\_,TE$ = $topk(E^{T}*(1-E^{A}),N-k)$   # Compute the index $TE$ of the selected text expert.
7:  Step 2. Expert specialization training for single modality
8:   $M[AE]$ = $M[AE](D^{A})$   # The audio expert undergoes specialized training on $D^{A}$.
9:   $M[TE]$ = $M[TE](D^{T})$   # The text expert undergoes specialized training on $D^{T}$.
10:  Stpe 3. Joint training of multi-modality experts
11:   $M^{\prime}$ = $M(D^{A}\cup D^{T})$
12:  return $M^{\prime}$, the MoE SLM model after modality expert partitioning and training.
3.3 Training Strategy for Modality Experts
In order to fully leverage the capabilities of both text and audio experts, it is necessary to conduct specialization training for the designated modality experts, enabling each group of modality experts to excel at processing inputs from their respective modalities. Finally, to handle multimodal information inputs simultaneously, cross-modal instruction data is used for joint training of the modality experts. The specific training process is illustrated in Figure 3.
Figure 3: The training process consists of three stages: stage 1 is for modality alignment, stage 2 is for training single modality experts, and stage 3 is for joint training of the modality experts.
Modality Aligment During the modality alignment stage, audio ASR data is used to align the semantic spaces of speech and text. In this stage, the main focus is on training the Adapter that connects the Audio Encoder and the LLM. Through modality alignment training, it is ensured that speech can be understood by the LLM in the same way as the meaning expressed by text.
Unimodal Expert Specialization Training The main purpose of unimodal specialization training is to enable modality experts to focus more on their own modality domains, so that they can fully utilize their expertise when collaborating with other multimodal experts. Since audio, like text, covers a wide range of domains, there are also multiple audio experts, each responsible for learning specialized knowledge in different areas, such as pitch, prosody, dialect, speech rate, and emotion. As shown in step 2 of Algorithm 1, audio and text experts are specialized and trained on their respective modalities. Since only unimodal specialization learning is required, the router is kept frozen to prevent it from becoming biased toward a single modality. During the specialization of audio experts, the router weight scores corresponding to text experts are set to zero. The remaining router weight scores are then normalized and redistributed among the audio experts. The same process is applied when specializing the text experts.
Joint Training of Modality Experts Multimodal joint training is mainly designed to enable the model to jointly train unimodal experts so that they can collaboratively process multimodal input information. As shown in Stage 3 of Figure 3, this stage uses cross-modal outputs from both modalities to jointly train the audio and text experts. Since all experts participate in training at this stage, the router is also unfrozen and trained, allowing it to better learn how to handle modality routing when processing multimodal joint inputs.
3.4 Audio Generation with Reinforcement Learning
To broaden the coverage of speakers and domains, it is inevitable that the pre-training data contains label noise and pronunciation errors, which can lead to hallucinations in the model. We observed that the quality of generated response audio is not consistent under different sampling parameters. To ensure more stable quality in response audio generation, we introduce a reinforcement learning stage to improve the stability of speech generation. Specifically, we use the text from the LibriSpeech panayotov2015librispeech set and have the model paraphrase it to obtain response audio and corresponding text. The response audio is then transcribed into recognized text using Whisper-large, and the word error rate (WER) is calculated as the reward score. All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) rafailov2023direct training as follows:
$\textit{L}_{DPO}(P_{\theta};P_{ref})=-\textit{E}_{(x,y_{w}^{A},y_{l}^{A})\sim D}[log\sigma(\beta log\frac{P_{\theta}(y_{w}^{A}|x))}{P_{ref}(y_{w}^{A}|x))}-\beta log\frac{P_{\theta}(y_{l}^{A}|x))}{P_{ref}(y_{l}^{A}|x))}],$ (7)
where $\theta$ is the policy model and $ref$ is the reference model. $x$ denotes the text input of the audio to be synthesized, $y_{w}^{A}$ denotes a good speech sequence, and $y_{l}^{A}$ denotes a bad speech sequence.
4 Experimental Results and Analysis
4.1 Experiment Settings
Models The overall architecture of the DeepTalk model is shown in Figure 2. We use the pre-trained DeepSeek-V2-Lite as the initial MoE-based LLM backbone, Whisper-medium as the audio encoder, and an MLP-based adapter to connect the audio encoder and the LLM. DeepTalk contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts. A sparse routing mechanism is adopted, where each token is processed by 6 experts. The model consists of 27 layers in total: the first layer is a dense layer, followed by 26 MoE layers. Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention dao2023flashattention with 16 attention heads. In stage 1, the learning rate is set to 2e-5. Since the original LLM has not undergone any pre-training on audio modalities, the learning rate is set to 1e-4 in stage 2.1 to accelerate convergence, 2e-5 in stage 2.2, and 5e-5 in stage 3. We use the AdamW loshchilov2017decoupled optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95. The final evaluation is conducted on LLM benchmarks, SQA, ASR, and TTS tasks.
Datasets In stage 1, WenetSpeech zhang2022wenetspeech is used for modality alignment. In stage 2.1, AudioQA-1M gao2025lucy is used for specialized training of audio experts. In stage 2.2, inhouse-60K, databricks-dolly-15k dolly2023introducing , MathInstruct yue2023mammoth (262K) and camel-ai-math li2023camel (50K) are used for specialized training of text experts. In stage 3, AudioQA-1M is used for joint training of modality experts. For the RL stage, we sampled 28K text entries from LibriSpeech panayotov2015librispeech to construct the audio preference pair dataset.
4.2 Evaluations
Evaluation on Text to Text
Figure 4: The modality load of the experts in the first layers.
The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/. As shown in Table 3, modular MLLMs cause minimal damage to the language abilities of the original text LLMs, primarily because the LLM is only responsible for generating text tokens. However, native MLLMs significantly impair the language capabilities of the original text LLMs, with an average relative loss of over 20% compared to their backbone LLMs. The DeepTalk model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
Evaluation on Speech to Text
Table 2: Performance on Spoken Question Answering
Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo chen2025minmo 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio huang2025step 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni wang2024freeze 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi defossez2024moshi 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice zeng2024glm 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY gao2025lucy 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend zhong2024moextend 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepTalk 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
The evaluation of speech-to-text is divided into two tasks: Spoken QA and ASR. As shown in the S $\rightarrow$ T column of Table 2, thanks to the preservation of the original LLM’s language capabilities by modular MLLMs, these models generally achieve leading performance. In contrast, native MLLMs cause significant degradation to the original LLM, resulting in increased hallucinations and thus lower scores. However, DeepTalk effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent. As a result, DeepTalk leads among native MLLMs. The ASR capabilities of the models are shown in Table 4, where DeepTalk, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
Table 3: Performance of Text to Text on LLM benchmark, * indicates that the average metric is calculated only on those test sets for which Qwen2.5-Omni has values. HS, WG, HEval denotes Hellaswag, Winogrande, HumanEval, respectively. “#AP” denotes number of activated parameters.
[…]
Table 4: Performance of Speech to Text on ASR. The metric for Chinese is CER, and for English WER. Lower values are better.
Model #A-Params WenetSpeech AIShell LibriSpeech
test-net test-meeting test test-clean test-other
Qwen2-Audio-base chu2024qwen2 7B 7.64 8.40 1.52 1.74 4.04
Baichuan-Audio-base li2025baichuan 7B 10.13 13.28 1.93 3.02 6.04
VITA fu2025vita 7B 12.15 16.53 - 8.14 18.41
Step-Audio-chat huang2025step 130B 9.47 10.83 2.14 3.19 10.67
Qwen2.5-Omni xu2025qwen2 7B 6.04 7.71 1.13 2.37 4.21
GLM-4-Voice zeng2024glm 9B - - 3.02 2.10 4.90
DeepTalk 2.4B 8.98 10.67 2.42 3.25 8.21
Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepTalk, where "topk" denotes the sampling parameter for audio tokens. Since the speech dialogue data used by DeepTalk contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English. After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in Recall@1, thereby making high-quality audio generation more stable. However, for more difficult tasks (test-hard), the supervised fine-tuning (SFT) model itself may not generate the correct result within the entire search space, so applying DPO brings little to no benefit.
Evaluation on Speech to Speech The evaluation of speech-to-speech is shown in the S $\rightarrow$ S column of Table 2. MoExtend refers to the results obtained by extending the original Deepseek-V2-Lite model with additional experts, adding six audio experts. However, since expanding the number of experts alters the distribution of the router, the performance is relatively suboptimal. DeepTalk, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts. This allows the audio experts to specialize while retaining the capabilities of the original text experts, thus achieving better performance among native MLLMs.
Latency We deployed DeepTalk on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module. The reported numbers are averaged over more than 10 samples. With a good network connection, the end-to-end latency of the entire system in half-duplex mode is 0.4362 seconds. The time cost for generating the first audio chunk is 0.3417 seconds. By subtracting the first chunk cost from the total latency, we estimate the network transmission cost to be 0.1062 seconds. The time cost for decoding a single token is 0.0205 seconds. All time measurements were conducted on a GPU with 80GB of memory.
5 Conclusion
To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepTalk, a MoE framework for adaptive modality expert learning. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs a performance loss lower than peer native MLLMs and on par with modular ones. Meanwhile, the end-to-end dialogue latency remains within 0.5 seconds, ensuring a seamless and intelligent speech interaction experience. To the best of our knowledge, this is the first MoE-based native multimodal large speech interaction model, which opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
Appendix A Dataset Details
As shown in Table 5, we provide the data and quantities used for training at each stage. Stage 1’s training data consists of speech-text pairs, which are used to align the semantic spaces of the speech and text modalities. AudioQA-1M contains conversational instruction data, with both input and output encompassing two modalities, aimed at fine-tuning the model to enhance its speech dialogue capabilities. Stage 2.2 utilizes pure text data to maintain the language proficiency of the text expert. During the RL stage, data from Librispeech is used to construct DPO preference pair data for speech generation.
Table 5: List of datasets used for each training stage.
Stage Dataset Size
1 WenetSpeech 10Kh
2.1 AudioQA-1M 1000K
2.2 In-House60k 60K
MathInstruct 262K
camel_ai_math 50K
databricks_dolly_15k 15K
3 AudioQA-1M 1000K
In-House60k 60K
MathInstruct 262K
camel_ai_math 50K
databricks_dolly_15k 15K
RL LibriSpeech-DPO 28K
Appendix B Experimental Setup Details
Table 6 outlines the main hyperparameters used during each stage of training. All ablation studies were conducted using a total batch size of 128, with learning rates of 1e-4, 2e-5, and 5e-5.
Table 6: Hyperparameters used in each stage training on DeepTalk.
Hyperparameter Stage 1 Stage 2.1 Stage 2.2 Stage 3 Stage RL
Learning rate 2e-5 1e-4 2e-5 5e-5 2e-5
LR schedule Cosine Cosine Cosine Cosine Cosine
Batchsize per GPU 16 8 8 8 8
Zero Zero2 Zero2-offload Zero2 Zero2-offload Zero3
Optimizer AdamW AdamW AdamW AdamW AdamW
Appendix C Ablation Study
C.1 Audio Experts Number
We conducted ablation experiments with varying numbers of audio experts, as shown in Table 7. The results demonstrate that as the number of audio experts increases, the language ability tends to be compromised, whereas the speech dialogue capability is enhanced. However, once the number of audio experts surpasses 24, both speech dialogue and language abilities begin to decline. This suggests that at this point, too many resources are allocated to audio experts at the expense of the text experts of the original text LLM, leading to a decrease in language ability. Moreover, the speech dialogue capability also falls after reaching a bottleneck concerning language ability.
Table 7: Ablation study on the impact of the number of audio experts on the capabilities of the two modalities.
Experts Num LLaMA Question Web Question TriviaQA Avg
T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S T $\rightarrow$ T S $\rightarrow$ S
6 69.0 59.7 46.6 23.1 58.2 27.5 57.9 36.8
12 61.0 60.0 39.2 26.7 49.7 29.1 50.0 38.6
18 59.0 59.0 32.1 28.3 42.6 33.7 44.6 40.3
24 48.0 48.0 23.7 22.9 31.9 30.2 34.5 33.7
C.2 The Modality Load of Different Layer Experts
Figure 5: The modality load of the experts in the middle layers.
Figure 6: The modality load of the experts in the last layers.
The expert load across different layers is illustrated in Figures 5 and 6, representing the modality load of an intermediate MoE layer and the final MoE layer, respectively. It is evident that there are significant differences in modality load across different layers. This indicates that the indices of audio experts in each layer are not the same.
C.3 Modality-Specific MoE
Table 8: Performance on Spoken Question Answering
Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
As shown in Table 8, we compared two settings: PureMoE without specifying modality experts and Random Modality-Specific MoE with randomly assigned experts. We found that if modality experts are not set, the model can only rely on self-learning the modality boundaries. During the early stages of training, the model does not clearly understand these boundaries, and the inclusion of audio data may significantly harm important text experts, leading to poorer final performance. The random assignment of modality experts was attempted three times, but results were inconsistent. If suitable experts were chosen as audio experts, the performance could be quite good, as seen in the second and third instances. However, if some crucial text experts were chosen as audio experts, the initial LLM suffers substantial damage, resulting in noticeable performance degradation. Using an adaptive modality assignment method effectively selects which experts are suitable as audio experts and text experts, achieving better performance.
Appendix D Border Impact
DeepTalk leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs. These limitations include the potential to generate inaccurate information or biased outputs. To address these issues, we enhanced the model’s speech interaction capabilities using AudioQA and performed speech-language instruction tuning on high-quality datasets. Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepTalk model in any downstream applications.